A 0.62-kg Slinky has 185 coils. When you and your friend stretch the Slinky by exerting a 4.0-N force on its ends, the separation between the adjacent coils is 5.0 cm.

A. Determine the speed of the transverse wave on the Slinky.
Answer #1

Length, L = (185 - 1) x 0.05 = 9.2 m

u = 0.62/9.2 = 0.0674 kg/m

Velocity, v = sqrt(T/u) = sqrt(4/0.0674) = 7.70 m/s

Comment in case any doubt please rate my answer ....